Hot water cools from 60°c to50°c in first 10 minutes and to 42°c in the next 10minutes the temperature of the surrounding is

Grade:11

1 Answers

Arun
6 years ago
Dear Rohit
 
average temp. in first case = (60+50) /2 = 55 C
average temp. in second case = (50+42) /2 = 46 C
as frm law of cooling
dT/dt = -k(T-To )  where To is temp. of surrounding
case 1 : dT = 10 C and dt = 10 mins
so,  1=-k(55-To )         ....(1)
case 2 : dT = 8 C and dt = 10 mins
so,  0.8=-k(46-To )       .........(2)
From 1 and 2 , we have
Temp of surr. as 10 C

Regards
run (askIITians forum expert)